OpenStudy (anonymous):

A price p (in dollars) and demand x for a product are related by 2x2+3xp+50p2=28000. If the price is increasing at a rate of 2 dollars per month when the price is 20 dollars, find the rate of change of the demand. Rate of change of demand =

OpenStudy (anonymous):

are the x an p values squared?

OpenStudy (anonymous):

is it 2x^2+3xp+5p^2=28000?

OpenStudy (anonymous):

Please reply so that I can know to continue.

OpenStudy (anonymous):

Anyways, So p=20 dp/dt=2 then find what you want to know and we want to know what dx/dt is, so differentiate the equation with respect to d/dt so we will get:

OpenStudy (anonymous):

Sorry the equation is like 2x^2+3xp+50p^2=28000

OpenStudy (anonymous):

2x dx/dt + 3x dp/dt +p dx/dt + 100p dp/dt = 0

OpenStudy (anonymous):

oo i think I know where I went wrong.. I didn't do 3p dx/dt but plz continue as I have tried this question 3x and have diff results which are all wrong T.T

OpenStudy (anonymous):

Also is it -3xp or +3xp?

OpenStudy (anonymous):

+3xp and shouldn't 2x dx/dt be 4x dx/dt?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Let me start over

OpenStudy (anonymous):

p and x are functions of t, so step 1 take the derivative of both sides

OpenStudy (anonymous):

d/dt(2x^2+3xp+50p^2)=d/dt(28000)

OpenStudy (anonymous):

You get 4xx' +3x'p+3xp'+100pp'=0

OpenStudy (anonymous):

now factor our x'

OpenStudy (anonymous):

Are you following so far?

OpenStudy (anonymous):

yeah I got that far

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

(4x+3p)x'+(3x+100p)p'=0

OpenStudy (anonymous):

then isolate right?

OpenStudy (anonymous):

(4x+3p)x'=-(3x+100p)p'

OpenStudy (anonymous):

x'=-(3x+100p)p'/(4x+3p)

OpenStudy (anonymous):

That would be part 1

OpenStudy (anonymous):

ok yeah this is exactly what I did - it ends up getting x=80 but plz continue

OpenStudy (anonymous):

now part 2. p=$20,

OpenStudy (anonymous):

2x^2+3x(20)+50(20)^2=28000.

OpenStudy (anonymous):

2x^2+60x+20000=28000

OpenStudy (anonymous):

2x^2+6x-8000=0

OpenStudy (anonymous):

x^2+3x-4000=0

OpenStudy (anonymous):

Now factor

OpenStudy (anonymous):

60 goes to 30

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

it should be x^2+30x-4000

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

nice catch

OpenStudy (anonymous):

and yeah I screwed up a sign in that equation maybe that's where I went wrong and tyty

OpenStudy (anonymous):

(x-50)(x+80)=0

OpenStudy (anonymous):

So x can be -80 and 50

OpenStudy (anonymous):

but it must be positve so its 50

OpenStudy (anonymous):

So now you know p=$20, p'=$2 per month, x =50 units

OpenStudy (anonymous):

now plug this into the previous x'

OpenStudy (anonymous):

x'=-(3(50)+100(20))(2)/(4(50)+3(20))

OpenStudy (anonymous):

solve it and it would give you the units per month

OpenStudy (anonymous):

omg kk thank you so much!!! I got it right! So how do I give you a good testimonial and all? hahahah

OpenStudy (anonymous):

Um, I guess you could do it if you click on my name. But I think you have to become a fan first. Something like that

OpenStudy (anonymous):

Ok became a fan! thanks again!

OpenStudy (anonymous):

yep, glad to help